### v1

@@ Appendix A The evaluations of clustering
[…]
@@
Calinski-Harabasz Index (CHI), also called Variance Ratio Criterion (VRC), measures clustering effectiveness using the ratio of the distance between groups and the distance within group. Larger value of CHI indicates better result of clustering. Assuming there are $k$ clusters in the clustering of variable $\mathbf{x}$, with $n_{i}$ samples in each cluster, we note the centroid of cluster $C_{i}$ as $\mathbf{c_{i}}$ and the overall centroid of the data as $\mathbf{c}$. The CHI can be expressed as
$\text{CHI}=\frac{\sum\limits_{i=1}^{k}n_{i}\|\mathbf{c}_{i}-\mathbf{c}\|^{2}/(k-1)}{\sum\limits_{i=1}^{k}\sum\limits_{\mathbf{x}\in C_{i}}\|\mathbf{x}-\mathbf{c}_{i}\|^{2}/(n-k)}.$ (19)
Appendix B Additional numerical results
We show the comparison of different estimations based on all 5 data-generating process in Table 4 below, where the results for DGP1 and DGP5 are also shown also in Table 2 in Section 3.2.
Table 4: Evaluations based on clustering results of estimated parameters $\hat{\bm{\beta}}$ by different methods
Evaluation Dimension Method RI ARI SC CHI(×$\times 10^{4}$)
DGP1 $\beta_{1}$ GWR 0.6411 0.2453 0.4537 0.2041
SCC 0.9117 0.7898 0.8685 2.3816
FLAT 0.9180 0.8051 0.8791 3.3122
$\beta_{2}$ GWR 0.7446 0.2769 0.4163 0.5611
SCC 0.9185 0.7490 0.8378 3.1291
FLAT 0.9289 0.7804 0.8577 4.0711
$\beta_{3}$ GWR 0.8084 0.3576 0.5163 0.9952
SCC 0.9346 0.7709 0.8713 4.3364
FLAT 0.9370 0.7818 0.8835 4.7318
DGP2 $\beta_{1}$ GWR 0.6220 0.1843 0.3160 0.2232
SCC 0.8816 0.7280 0.8373 1.5141
FLAT 0.8932 0.7519 0.8492 2.0688
$\beta_{2}$ GWR 0.7210 0.2173 0.3331 0.5846
SCC 0.8942 0.6826 0.8137 2.4471
FLAT 0.9074 0.7176 0.8307 3.2222
$\beta_{3}$ GWR 0.7800 0.2933 0.4142 1.0047
SCC 0.9002 0.6753 0.8432 2.9139
FLAT 0.9107 0.7056 0.8441 3.6620
DGP3 $\beta_{1}$ GWR 0.6298 0.1702 0.2606 0.2975
SCC 0.8525 0.6607 0.7910 1.2801
FLAT 0.8697 0.6946 0.8236 1.8807
$\beta_{2}$ GWR 0.7027 0.1853 0.2720 0.6311
SCC 0.8732 0.6185 0.7928 2.0624
FLAT 0.8904 0.6672 0.8204 2.6647
$\beta_{3}$ GWR 0.7619 0.2607 0.3454 1.0801
SCC 0.8832 0.6296 0.8217 2.4289
FLAT 0.8959 0.6589 0.8281 2.8204
DGP4 $\beta_{1}$ GWR 0.6206 0.1481 0.2246 0.3315
SCC 0.8286 0.6071 0.7702 1.0753
FLAT 0.8516 0.6545 0.7990 1.5322
$\beta_{2}$ GWR 0.6968 0.1681 0.2144 0.7392
SCC 0.8559 0.5701 0.7715 1.8524
FLAT 0.8674 0.6030 0.7841 2.5714
$\beta_{3}$ GWR 0.7467 0.2370 0.2862 1.1285
SCC 0.8661 0.5871 0.7931 2.0760
FLAT 0.8799 0.6150 0.8172 2.4064
DGP5 $\beta_{1}$ GWR 0.6101 0.1260 0.1879 0.3858
SCC 0.8056 0.5650 0.7494 0.7449
FLAT 0.8227 0.5951 0.7680 1.0421
$\beta_{2}$ GWR 0.6760 0.1487 0.1363 0.7457
SCC 0.8307 0.5023 0.7572 1.6862
FLAT 0.8471 0.5386 0.7754 2.3999
$\beta_{3}$ GWR 0.7176 0.2061 0.2091 1.3860
SCC 0.8640 0.5660 0.7871 2.0479
FLAT 0.8718 0.5892 0.8066 2.1975
Appendix C Details of the real-data analysis

### v2

@@ 5.2 The evaluations of clustering
Since the RI does not yield a value of zero for random clustering results, the Adjusted Rand Index (ARI) is frequently employed. It corrects for the influence of random clustering, with a value range from -1 to 1, where an ARI of 0 indicates purely random clustering. The ARI addresses the limitation of RI, which tends to exhibit spuriously high values when the number of clusters is large due to chance agreements. Additionally, it enables direct comparison of results across different clustering algorithms. The expected value of RI under random clustering is denoted as $\text{E}(\text{RI})$. Then, the ARI is computed as
@@
Calinski-Harabasz Index (CHI), also called Variance Ratio Criterion (VRC), measures clustering effectiveness using the ratio of the distance between groups and the distance within group. Larger value of CHI indicates better result of clustering. Assuming there are $k$ clusters in the clustering of variable $\bm{x}$, with $n_{i}$ samples in each cluster, we note the centroid of cluster $C_{i}$ as $\mathbf{c_{i}}$ and the overall centroid of the data as $\mathbf{c}$. The CHI can be expressed as
$\text{CHI}=\frac{\sum\limits_{i=1}^{k}n_{i}\|\mathbf{c}_{i}-\mathbf{c}\|^{2}/(k-1)}{\sum\limits_{i=1}^{k}\sum\limits_{\bm{x}\in C_{i}}\|\bm{x}-\mathbf{c}_{i}\|^{2}/(n-k)}.$ (19)
5.3 Details of the real-data analysis